A drop of water fell on the surface of a still water,so a 120 wave ripples are produced during half minute.If the radius of external circle is 3m, the speed of prpagation of wave equals.....?
1
Expert's answer
2020-01-20T05:18:59-0500

Distance between waves:


"d=\frac{3}{120}=0.025\ m=2.5\ cm"

Ths speed of propagation of wave equals


"v=\frac{d}{t}=\frac{2.5}{0.5}=5\frac{cm}{min}"